This is the question from #1:

The standard mixture you were given was prepared by mixing 18.0 mL of 0.200 M Fe(NO3)3 with 2.0 mL of 0.00200 M KSCN. Assuming that all of the SCN- reacted in this mixture, what is the final concentration of FeSCN2+ in the mixture?

For that, I got 2*10^-4M as the answer.

I need help with the second problem:

Use the value from #1 and the absorbance of the solution (recall A = 2-log %T) to determine the [FeSCN2+] in the other solutions based on their absorbances. [hint: Beer's Law says that absorbance is directly proportional to concentration]

Absorbances are as follows:

a) 2.304

b) 2.684

c) 2.41

d) 2.383

e) 2.343

f) 2.381
Answer #1

equations needed:
[SCN–]eq = [SCN–]i – [FeNCS2+]eq
[FeNCS2+]eq = (Aeq/Astd) * [FeNCS2+]std
let,
Aeq is given= 2.4
Astd is given= 2.68
[FeNCS2+]std = .0002

plug it in!

[FeNCS2+]eq = (2.4/2.68) * .0002 = 1*10^-3 M